# Gaga: Group Any Gaussians via 3D-aware Memory Bank

### v2

@@ 4.3 Further Implementation Details
5 Supplementary Ablation Study
5.1 Ablation Study on Overlap Threshold
We conduct additional ablation studies on the Gaussian overlap threshold using the Replica dataset [27], with SAM [16] as the 2D segmentation model. During the group ID assignment process, if none of the existing groups in the memory bank exceed the overlap threshold with the current mask, we add the mask as a new group, indicating the discovery of a new 3D object.
@@ 5.1 Ablation Study on Overlap Threshold
5.2 Comparison on Mask Association Methods
Figure 13: Visual comparison between different mask association methods. Our 3D-aware memory bank offers more detailed associated masks, accurately tracks identical objects in the scene and assigns them different mask IDs. Conversely, video object tracker leaves empty regions in positions where it cannot track masks, and it struggles to provide consistent masks for the same object across views.
We present visual comparison results for two mask association methods, video tracker [7] utilized by [32] and Gaga’s 3D-aware memory bank, in Fig. 13. In the “garden” scene of the MipNeRF 360 dataset, the video tracker struggles to track objects in the background, whereas Gaga provides associated results for each mask. For the scene in the ScanNet dataset, the video tracker fails to distinguish between four identical sofas, resulting in multiple masks for the same object. Additionally, it assigns different mask IDs to the table in two views. In contrast, Gaga precisely locates each object, leading to improved mask association results and better pseudo labels for training segmentation features.
6 Limitations

### v3

@@ C.3 Further Implementation Details
Appendix D Supplementary Ablation Study
D.1 Ablation Study on Overlap Threshold
We conduct additional ablation studies on the Gaussian overlap threshold using the Replica dataset replica , with SAM sam as the 2D segmentation model. During the group ID assignment process, if none of the existing groups in the memory bank exceed the overlap threshold with the current mask, we add the mask as a new group, indicating the discovery of a new 3D object.
@@ D.1 Ablation Study on Overlap Threshold
D.2 Comparison on Mask Association Methods
Figure 13: Visual comparison between different mask association methods. Gaga offers more detailed associated masks, accurately tracks identical objects in the scene and assigns them different mask IDs. Conversely, Gaussian Grouping leaves empty regions in positions where it cannot track masks, and it struggles to provide consistent masks for the same object across views.
We present visual comparison results for two mask association methods, video tracker deva utilized by gg and Gaga’s 3D-aware memory bank, in Fig. 13. In the “garden" scene of the MipNeRF 360 dataset, the video tracker struggles to track objects in the background, whereas Gaga provides associated results for each mask. For the scene in the ScanNet dataset, the video tracker fails to distinguish between four identical sofas, resulting in multiple masks for the same object. Additionally, it assigns different mask IDs to the table in two views. In contrast, Gaga precisely locates each object, leading to improved mask association results and better pseudo labels for training segmentation features.
Appendix E Limitations